If you toss 4 coins what is the probability of getting all heads?
Solution to question is P(H) and P(H) and P(H) and P(H), which is 1/2 * 1/2 * 1/2 * 1/2 = 1/16 = 0.0625.
4/27
What is the probability of rolling a number that is a multiple of 3 on a die?
There are two multiples of three on a die and there are six possible outcomes.
So, the probability is 2/6 = 1/3 = 0.3333 or 33.33%
What is the probability of rolling a number less that 3 on a number cube?
If it is fair die, then the probability is 1/3.
The best known is British Standards Institute, but there are at least 50 other meanings to the abbreviation.
How many different outcomes would be possible if you rolled a single die and then two coins?
The single die can land in 6 different ways. For each of those . . .
The first coin can land in 2 different ways. For each of those . . .
The second coin can land in 2 different ways.
Total number of different possible outcomes = (6 x 2 x 2) = 24
IF the coins have different colors so that you can tell them apart.
If you can't tell the coins apart, then there are only 12 different outcomes.
How many outcomes are there for tossing an odd number on a number cube?
there are three distinct outcomes1,3or5. But there is no probability which will come.
What is the probability of rolling 7 or 11 with 2 dice?
The probability of rolling a 7 on a pair of 2 dices is: P(7) = 6/36.
The probability of rolling a 11 on a pair of 2 dices is: P(11) = 2/36.
The probability of rolling 7 or 11 is the sum of both single probabilities:
P(7 or 11) = 6/36 + 2/36 = 8/36 = 2/9 = 0.2222... ≈ 22.2%.
Explanation.
When rolling a pair of fair dice there are 36 possible equiprobable events or
outcomes. They are:
(1,1), (1,2), (1,3), (1,4), (1,5), (1,6),
(2,1), (2,2), (2,3), (2,4), (2,5), (2,6),
(3,1), . . . . . . . . . . . . . . . . . , (3,6),
(4,1), . . . . . . . . . . . . . . . . . , (4,6),
(5,1), . . . . . . . . . . . . . . . . . , (5,6),
(6,1), . . . . . . . . . . . . . . . . . , (6,6).
Each of these 36 events has the probability of happening of 1/36.
For the case of the sum of both dice to be 7, there are 6 possible outcomes, so the
probability of the event sum equals 7 is: P(7) = 6/36 = 1/6.
For the case of the sum of both dice to be 11, there are 2 possible outcomes, (6,5)
and (5,6), so the probability of the event sum equals 11 is: P(11) = 2/36 = 1/18.
In the year 1964 there was a US Presidential election between the Democratic candidate Lyndon Johnson and the Republican Barry Goldwater, in which Johnson won. This was also a year in which the Beatles gained increasing popularity around the world. There were many other notable events although those are the two that I particularly remember.
What is the probability that a month picked at roandom ends with the later y?
If using the months' names in the English language, it is 1/3.
What are the chances of a female becoming a top executive?
Theorically, women have the same potentiality as men, I would even dare to change 'same' with 'more', since we are usuallymore practical, fast and astute... The reason why there are few top executive women is not to be looked there, but mostly in 'Time Management'. Having such a career requires long hours, days and weeks far from home - and having a family in such circumstances is not only difficult, it's impossible for most of us. How many men would like to stay home taking care of their children and house while their wives are abroad meeting (almost surely) other men? ;) In any case, I know it sounds like a cliché, but really most business men are strongly backed (advised, supported) by their wives - women may not have a highlighted place in the business world, but they are there, working in the backstage.
What has the author H H T Cleife written?
H. H. T. Cleife has written:
'England's greatest national sin'
It is (the number of sectors which are numbered 2) divided by 6
To work this out, you first need to calculate the probability of the first 8 passing, and the final 2 failing. That is 97%8x3%2 which equals 0.070536902349392649%
The next step is to find out how many different ways the people can be ordered. We have 10 in total and want to order 8, so the number is 10C8, which is 45. Multiply this by the above probability and you get 3.174160605722669205%
Thus the probability that 8 of the next 10 will pass is 3.17%
It is (5/14)^2 = 0.1276, approx.
There are 4 aces, so the odds of drawing an ace are 4/52, or 1/13.
Who introduce complementary events?
It was quite likely to be a prehistoric event since all events are dichotomous: either they happen or they don't happen. Either the sabre tooth will kill me or it won't, either I will find enough prey or I won't. There is no other outcome. So, in each case, the pair of events are mutually complementary.
What is the probability of getting an eight in a deck of cards?
If you're including jokers, 4/52-if you're not including jokers-4/50 (8%).